### v2

# In-context KV-Cache Eviction for LLMs via Attention-Gate

Abstract
The KV-Cache technique has become the standard for the inference of large language models (LLMs). It caches states of self-attention to avoid recomputation. Yet, it is widely criticized that KV-Cache can become a bottleneck of the LLM inference system, especially when confronted with ultra-large models and long-context queries. A natural remedy is to discard the KV-Cache for less important tokens, with StreamingLLM (Xiao et al., 2024) as an example, but the used static eviction strategies cannot flexibly adapt to varying contexts. Remedies like H2O (Zhang et al., 2024) leverage accumulative attention scores to perform dynamic eviction but suffer from the attention bias issue in capturing contextual information. This paper bridges this gap by devising a parameterized KV-Cache eviction mechanism, dubbed as Attention-Gate, which accepts the whole context as input and yields eviction flags for each token to realize in-context eviction. The subsequent self-attention module proceeds according to the flags and only the KV states for the remaining tokens need to be cached. The Attention-Gates can vary among different heads and layers and be trivially plugged into pre-trained LLMs, tuned by cost-effective continual pre-training or supervised fine-tuning objectives to acquire what to discard. The computational and memory overhead introduced by Attention-Gates is minimal. Our method is validated across multiple tasks, demonstrating both efficiency and adaptability. After a highly efficient continual pre-training, it achieves higher average accuracy and evicts more tokens compared to traditional training-free methods. In supervised fine-tuning, it not only evicts many tokens but also outperforms LoRA-finetuned LLMs on some datasets, such as RTE, where it improves accuracy by 13.9% while evicting 62.8% of tokens, showing that effective eviction of redundant tokens can even enhance performance.
1 Introduction
Large language models (LLMs) (Dubey et al., 2024; Team et al., 2024; Chiang et al., 2023) have achieved remarkable success across a wide range of natural language processing tasks. A key technique that has enabled efficient LLM inference is KV-Cache, which stores transient attention keys and values to avoid recomputation. However, as the size of LLMs continues to increase and the demand for handling long-context queries grows, the KV-Cache has emerged as a significant bottleneck. Storing attention states for numerous tokens can lead to considerable memory overhead and data transfer among the memory hierarchy results in substantially increased inference time.
Studies have shown that sparsity is a natural phenomenon in attention mechanisms, with many tokens being redundant for inference (Zhang et al., 2024). This suggests that retaining all tokens in the KV-Cache is unnecessary. Existing works have explored this insight to compress KV-Cache using static strategies or hinging on accumulative attention scores. StreamingLLM (Xiao et al., 2024) is a representative of the former by retaining a fixed window of beginning and recent tokens in the KV-Cache but it struggles to flexibly adapt to specific contexts. E.g., in sentiment analysis, retaining the token “cute” in “a cute cat” is crucial, while in object recognition, the token “cat” would be more important. H2O (Zhang et al., 2024), on the other hand, employs a token-adaptive approach, using local accumulative attention scores to determine which tokens to evict. However, it is criticized that in practice, H2O suffers from the attention bias issue (Oren et al., 2024), with a tendency to over-prioritize either the initial or recent tokens.
Figure 1: KV-Cache eviction patterns across different layers and attention-heads, visualized for 4 samples from the PIQA dataset (top row) and 4 samples from the BoolQ dataset (bottom row), using AG fine-tuned Llama2-7B models. Black areas represent tokens that are neither computed nor stored in the KV-Cache. The variability of eviction patterns across tasks, prompts, layers, and attention-heads demonstrates the dynamic nature of our method. A common trend observed is that deeper layers tend to mask more KV-Cache states, with some in deeper layers being entirely masked.
To overcome these challenges, we introduce a parameterized KV-Cache eviction mechanism named the Attention-Gate (AG), to perform reliable in-context eviction. AG is positioned before a self-attention layer within the model. It takes a sequence of token features as input and generates eviction flags for the tokens, indicating whether a token should be excluded from subsequent self-attention computations. Tokens that are evicted do not require their KV states to be cached. AGs can be seamlessly integrated into pre-trained LLMs and tuned by minimizing the language modeling loss. Ideally, AGs can automatically learn to discern the most relevant tokens for the current context without manual intervention. In practice, we can implement the AG as a self-attention layer with much fewer heads than the original model (e.g., 4 v.s. 32). This way, the parallel computational capabilities of the hardware can be harnessed to minimize the extra overhead introduced by AGs.
AG is empirically shown to enjoy high training efficiency, e.g., only four NVIDIA 4090 GPUs and a dataset of 5,000 samples are required for continual pre-training when applying AGs to LLaMA2-7B (Touvron et al., 2023). This alleviates concerns about the computational overhead related to trainable eviction strategies (Zhang et al., 2024; Chen et al., 2024) and amplifies the performance superiority of our approach over existing training-free approaches. As illustrated in Figure 1, AG generates different eviction strategies across different layers and attention-heads for different tokens, demonstrating its adaptability to the diverse requirements of each component in the model.
To validate the effectiveness of our method, we conduct extensive experiments across multiple benchmarks. After efficient continual pre-training (CPT), our approach outperforms traditional training-free eviction strategies, such as StreamingLLM and H2O, in both accuracy and token eviction rates. In supervised fine-tuning (SFT), our method not only evicts a significant number of redundant tokens but also maintains or surpasses the performance of LoRA-finetuned LLMs. For example, on the RTE dataset, our approach improves accuracy by 13.9% while evicting 62.8% of tokens, demonstrating that selective token eviction can enhance performance. In summary, the Attention-Gate mechanism provides a scalable and efficient solution for KV-Cache management, addressing the limitations of traditional training-free methods.
Figure 2: An overview of Attention-Gate (AG) for KV-Cache eviction. AG is a lightweight learnable module placed before each MHA layer. Given the input hidden states, it determines for each head whether to retain or discard the key and value tokens in the KV-Cache. In the attention weights, this corresponds to masking out columns for the evicted keys, while keeping the diagonal intact to ensure the query interacts with its own key.
2 Related Work
As large language models (LLMs) scale in both size and input sequence length, optimizing their efficiency has become increasingly important, particularly in addressing space and time complexity. A significant bottleneck lies in the attention mechanism, which demands considerable computational and memory resources, especially for long sequences.
Computation and memory challenges. The computational cost of attention is driven by its quadratic complexity, while the memory burden stems from the need to store the KV-Cache during inference to avoid recomputation. Although several works have focused on reducing the computational load of attention mechanisms, many fail to address the memory constraints that remain a critical issue. Techniques like Reformer (Kitaev et al., 2020) and FlashAttention (Dao et al., 2022; Dao, 2023) mitigate memory usage for long sequences but still require substantial cache storage. On the other hand, approaches such as MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023) prioritize memory compression without tackling computational complexity. These solutions tend to focus on either memory or computational efficiency, rarely addressing both simultaneously.
KV-Cache eviction strategies. To address both memory and computational challenges, KV-Cache eviction has emerged as an effective strategy. Existing approaches can be categorized into static and adaptive methods based on accumulative attention scores.
Static strategies, such as those used in Sparse Transformers (Child et al., 2019), employ fixed pruning patterns, such as Strided and Fixed Attention. While effective in some cases, these approaches are not adaptive to specific contexts, often sacrificing accuracy. StreamingLLM (Xiao et al., 2024) tackles the Attention Sink phenomenon, where attention scores concentrate on initial tokens, by retaining these tokens along with a fixed window of recent tokens. While this improves performance, static approaches generally lack the flexibility needed to adapt to different tokens, attention-heads, or layers.
Strategies using accumulative attention scores offer more flexibility by dynamically identifying important tokens. For instance, SpAtten (Wang et al., 2021) employs Accumulative Attention Scores (A2S), which sum the softmax outputs for each token to measure its importance. This approach allows selective token pruning in subsequent layers, effectively reducing computational complexity without the need for retraining. H2O (Zhang et al., 2024) extends this concept to decoder-based models, using local A2S statistics for adaptive eviction in autoregressive generation. However, H2O suffers from the attention bias issue (Oren et al., 2024), particularly in long-context inputs. Several follow-up works have aimed to address this limitation. NACL (Chen et al., 2024) introduces random eviction to mitigate attention bias, while A2SF (Jo & Shin, 2024) incorporates a Forgetting Factor. However, none of these approaches fully resolve the underlying problem.
More Adaptive strategies. Although strategies based on accumulative attention scores provide more flexibility than static methods, they still have notable limitations. For instance, H2O (Zhang et al., 2024) applies the same token eviction ratio across all attention heads, restricting the adaptability of the method. FastGen (Ge et al., 2023), on the other hand, introduces a different approach by hybridizing KV-Cache compression policies and applying adaptive strategies to each attention head. However, it focuses on the decoding stage and neglects the importance of the prefilling stage. Learnable eviction strategies, on the other hand, offer greater flexibility by enabling different layers and attention heads to adopt heterogeneous eviction policies. However, such strategies have been relatively underexplored, likely due to concerns about the computational overhead they may introduce (Zhang et al., 2024; Chen et al., 2024). Nonetheless, task-specific training is essential for optimizing performance across different contexts. For example, a recent approach (Anagnostidis et al., 2024) introduces a learnable mechanism for dropping uninformative tokens, but it faces difficulties in batched generation and does not account for continual pre-training or decoding-only LLMs. Despite these challenges, learnable strategies have strong potential to improve performance across a variety of tasks by allowing models to adapt their eviction strategies to meet task-specific requirements.
3 Method
This section first briefly reveals the basics of multi-head attention and KV-Cache and then describes the proposed Attention-Gate (AG) mechanism for in-context KV-Cache eviction for LLM inference acceleration. An illustrative overview of AG is presented in Figure 2.
3.1 Preliminary
Multi-Head Attention (MHA) (Vaswani et al., 2017) is a core component of the Transformer architecture, widely used in most LLMs. MHA enables the model to capture dependencies across different tokens in a sequence by computing self-attention across multiple attention-heads. Each head attends to different parts of the input sequence independently, allowing the model to capture various aspects of the relationships between tokens. The outputs of these attention heads are then concatenated and projected through a final output matrix.
KV-Cache is employed to store the key and value representations of tokens from previous time steps during inference. This prevents redundant recomputation of these representations for every new token generated, significantly speeding up auto-regressive generation. The inference process in auto-regressive transformers, such as Llama (Touvron et al., 2023), can be divided into two stages: prefilling and decoding. 1. In the prefilling stage, the model processes the entire input sequence, generating key-value pairs for each attention-head in each layer. These key-value pairs are stored in the KV-Cache for future reuse. 2. In the decoding stage, when generating the next token, the model only computes the key and value for that new token and appends them to the cached key-value pairs. This reduces the computational load, as the previously cached key-value pairs do not need to be recomputed. The process continues token by token until the sequence generation is complete.
KV-Cache plays a critical role in improving the efficiency of LLM inference, especially in scenarios where long sequences are processed. However, the size of the KV-Cache grows with the input sequence length, leading to substantial memory overhead. Efficiently managing this cache while maintaining model performance has become a key challenge in scaling LLMs to longer contexts.
3.2 Limitations of Traditional Eviction Strategies
Lack of flexibility. Flexibility in KV-Cache eviction strategies spans across various dimensions, including token-specific, attention-head-specific, layer-specific, task-specific, and model-specific adaptability. Static eviction policies offer no such flexibility, requiring manual intervention to adjust them. While methods based on accumulative attention scores (e.g., H2O (Zhang et al., 2024)) have improved upon static strategies by introducing token-level and head-level adaptability, they are still limited. Specifically, in H2O, the eviction ratio remains uniform across all attention heads, restricting head-level flexibility. Flexibility is crucial in eviction strategies because models process varying types of data across different contexts, requiring the ability to selectively retain important tokens while discarding others. Without this adaptability, the model may inefficiently store redundant information, increasing memory usage and slowing down inference times.
Absence of global statistics. Eviction strategies that rely on accumulative attention scores, such as H2O, base their calculations on local statistics, which often results in attention bias issues (Oren et al., 2024). This local approach may misjudge a token’s importance, especially when the context spans long sequences. While several methods (Chen et al., 2024; Jo & Shin, 2024) have proposed solutions to mitigate this bias, they are still fundamentally rooted in local statistics, addressing the issue only partially. A more effective solution would incorporate global statistics to ensure that the eviction process is based on a more comprehensive understanding of the context, thereby reducing biases and improving the accuracy of token retention.
Inefficiency. Efficiency is another challenge for traditional methods such as H2O and FastGen (Ge et al., 2023). These methods operates in a token-by-token manner for each decoding step, sequentially deciding which key-value pairs to evict from the cache. Additionally, H2O determines which tokens to evict after computing the attention scores, meaning some computation is wasted on tokens that are later discarded. This sequential approach hinders efficiency, especially when handling large batches of data or when processing long sequences.
Our Attention-Gate mechanism is designed to address the aforementioned limitations.
3.3 Attention-Gate
The Attention-Gate (AG) is a lightweight, trainable module designed to determine which tokens in the KV-Cache of each attention-head should be retained or discardedAG is positioned before the Multi-Head Attention (MHA) layer and operates by generating eviction flags for each token, guiding both the computation of attention scores and the management of the KV-Cache.
Input and output. AG takes the hidden states as input and generates binary flags as output, one for each token in every attention head. These flags indicate whether the corresponding token’s key-value pairs should be stored or discarded in the KV-Cache. Specifically, the input is a hidden state matrix $X\in\mathbb{R}^{n\times d}$, where $n$ represents the sequence length, and $d$ denotes the dimensionality of the hidden states. The output consists of independent binary decisions for the retention or eviction of each token’s key-value pairs across all attention heads in the MHA module.
Local or global? The structure of AG offers two choices: using either local or global information to guide the eviction policy. 1. Local information: The simplest approach leverages a linear layer where each token only considers its own hidden state without taking other tokens in the sequence into account. This method is computationally efficient and straightforward, making it appealing for its simplicity. 2. Global information: On the other hand, a more comprehensive approach adopts an attention-like structure, where tokens can aggregate information from across the entire sequence. This enables the eviction policy to reflect the global in-context information, making decisions that are informed by the broader context. Naturally, this method is computationally heavier than the local approach, but it offers a more accurate reflection of the sequence’s token importance.
In our design, we opted for the attention-like structure because global context is essential for making effective eviction decisions. As demonstrated in Section 4.4, using an attention-like structure significantly outperforms the linear layer, as the latter struggles to properly evict tokens. This outcome is expected, as determining whether a token is redundant or crucial for inference often requires a global understanding of the sequence. A local approach, like a linear layer focusing on individual tokens, lacks the broader perspective necessary to make accurate decisions about which tokens are important. This finding also underscores the limitations of methods like H2O, which rely on local statistics rather than global ones, as mentioned in Section 3.2.
Softmax or Sigmoid? After gathering information through the attention-like structure, the next step is to compute the eviction probability for each token. Common activation functions such as Softmax and Sigmoid can be used to generate these probabilities.
In our case, we choose the Sigmoid function for a key reason: while Softmax normalizes probabilities across the entire sequence, introducing competition between tokens, Sigmoid treats each token independently. This independence is crucial because we do not want a token’s likelihood of being evicted to depend on the eviction probabilities of other tokens in the sequence.
Eviction flags. After computing the eviction probabilities for each token, we apply a threshold $\tau$. If the probability of retaining a token exceeds $\tau$, that token is retained in the KV-Cache; otherwise, it is discarded. In the attention matrix, the columns corresponding to evicted tokens are masked out. However, the diagonal elements, where a token attends to itself, are always preserved. This guarantees that each token continues to interact with its own key.
Prefilling or decoding? AG is primarily applied during the prefilling stage, where the full sequence is available. By making eviction decisions before the MHA layers, AG effectively manages the KV-Cache during this phase. In contrast, AG is not used during the decoding stage to avoid adding complexity to the inference process. Using AG in decoding could slow down inference and increase the cache memory footprint, as additional keys and values from the attention-like structure would need to be stored.
In summary, AG dynamically determines which tokens should be retained, optimizing both KV-Cache usage and attention score computation by utilizing an attention-like structure, Sigmoid activation function, and a threshold-based decision for token eviction.
3.4 Training Implementation
To train the Attention-Gate (AG) effectively, this section outlines the key components of the training process. For more comprehensive details, please refer to Section 4 and Appendix A.
Eviction Loss. To encourage the eviction of unnecessary tokens, we introduce a dedicated loss function called the Eviction Loss. This loss encourages the model to discard as many tokens as possible, which is defined as:
$\ell_{\text{evict}}=\alpha\cdot\frac{1}{L}\sum_{l=1}^{L}\text{ReLU}\left(\overline{\text{AG}}_{l}-\beta\right),$ (1)
where $L$ is the number of layers, and $\overline{\text{AG}}_{l}$ represents the average output of the AG module in layer $l$. In this formula, $\alpha$ adjusts the intensity of KV-Cache eviction, while $\beta\in[0,\tau]$ ensures that eviction does not become overly aggressive.This loss function works alongside the auto-regressive loss to balance token eviction with maintaining model performance. The loss allows for adaptive eviction across layers and attention-heads. As shown in Figure 1, after training, the model learns to increase token eviction in deeper layers, which is not achievable with training-free methods like H2O (Zhang et al., 2024) or StreamingLLM (Xiao et al., 2024).
Initialization. We initialize the AG parameters using Xavier initialization (Glorot & Bengio, 2010) to provide a stable starting point for learning. Additionally, a small constant $\gamma\geq 0$ is added inside Sigmoid, ensuring that the initial retention probabilities are close to 1. This encourages the model to retain most tokens early in training, allowing it to learn which tokens are important more gradually.
Handling non-differentiability. Directly applying the threshold-based gating mechanism from Section 3.3 would lead to non-differentiable gradients during training due to the hard thresholding’s discrete nature. To resolve this, we employ the Straight-Through Estimator (STE) (Yin et al., 2019), which allows gradients to flow through discrete decisions by approximating them during the backward pass. Specifically, during backpropagation, instead of using the hard 0 or 1 values obtained from comparing against the threshold, we utilize the smooth output of the Sigmoid function. This approach ensures smooth gradients and enables effective training of the AG while preserving its binary behavior during the forward pass.
4 Experiments
This section consists of three main parts. First, we evaluate the performance of AG in two scenarios: continual pre-training (CPT) and supervised fine-tuning (SFT). For CPT, we compare AG with classic methods on performance and eviction rates (Table 2). For SFT, we benchmark AG against vanilla LoRA fine-tuning to highlight its task-specific adaptability (Section 4.2). Second, we provide visualization of selected examples to demonstrate the core characteristics of AG (Section 4.3). Finally, we conduct ablation studies to provide further insights into the effectiveness of AG (Section 4.4).
4.1 Continual Pre-training
Table 1: Performance comparison of Llama2-7B and various KV-Cache eviction strategies across seven tasks. The table reports accuracy (Acc.) for Llama2-7B and all eviction methods, with Llama2-7B serving as the upper bound for accuracy. Metric %Eviction refers to the mean KV-Cache eviction ratio, representing the percentage of tokens evicted from the KV-Cache. The eviction ratio is fixed at 50% for the baseline methods, including a local strategy (retaining only recent tokens), StreamingLLM, and H2O. In contrast, our method, which includes continual pre-training, achieves higher average accuracy while maintaining a higher average %Eviction.
Metric PIQA ARC-C ARC-E RTE COPA BoolQ HellaSwag Avg.
Llama2-7B Acc. 76.33 37.29 51.32 51.99 62.00 69.94 68.16 59.58
Local Acc. 69.97 31.86 48.68 51.99 60.00 57.86 37.08 51.06
StreamingLLM Acc. 72.69 33.22 51.15 50.18 63.00 62.05 40.85 53.31
H2O Acc. 75.9 33.22 52.03 52.71 47.00 67.37 66.32 56.36
Ours Acc. 76.17 33.90 49.03 52.35 63.00 67.52 66.33 58.33
%Eviction 54.29 51.03 51.05 46.70 40.02 57.75 52.16 51.87
Table 2: Performance of Llama2-7B with LoRA fine-tuning and our method on six downstream tasks. In addition to the LoRA fine-tuning targets, our method makes the AG modules learnable. Two settings for $\alpha$ (0.5 and 1) are tested. Our method maintains comparable or better accuracy while achieving a higher eviction ratio, demonstrating its task-specific adaptability in managing token eviction without significant accuracy loss.
Metric PIQA ARC-C RTE COPA BoolQ OBQA Avg.
Fine-tuned Llama2-7B Acc. 82.92 60.34 64.98 92.00 88.10 78.80 77.86
Ours $(\alpha=1)$ Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
%Eviction 66.16 48.31 65.47 45.40 67.46 67.17 60.00
Ours $(\alpha=0.5)$ Acc. 81.50 57.63 74.01 95.00 87.00 79.20 79.06
%Eviction 64.96 36.45 62.80 34.77 67.31 66.65 55.49
4.1.1 Setup
Models and datasets. We use Llama2-7B (Touvron et al., 2023) as our primary base model due to its strong performance and popularity within the AI community. Additionally, we validate the feasibility of our approach on Mistral-7B (Jiang et al., 2023), with results provided in Table 4. For continual pre-training, we select a subset of the Redpajama dataset (Computer, 2023) containing approximately 5,000 samples  [^1] We sampled 4,997 samples proportionally from each subset of RedPajama.. To assess the effectiveness of our method, we evaluate it on seven widely recognized benchmarks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), and HellaSwag (Zellers et al., 2019). All evaluations are conducted in a zero-shot setting, with accuracy assessed using OpenCompass (Contributors, 2023).
Training details. During continual pre-training (CPT), only the AG modules are learnable, while the rest of the model’s parameters are frozen. For the Llama2-7B model, the hyperparameters from Section 3.4 are used: $\tau=0.5$, $\gamma=2$, $\alpha=5$, and $\beta=0.4$. The model was trained for one epoch.
Metrics and Baselines. The metrics considered include accuracy and the mean eviction ratio for all KV-Cache. The accuracy of vanilla Llama2-7B on these datasets is used as the upper bound, with the goal of minimizing accuracy loss while maximizing the mean eviction ratio. For KV-Cache eviction strategies, we select a local method (retaining only recent tokens) and StreamingLLM (Xiao et al., 2024) as representatives of static strategies, while H2O (Zhang et al., 2024) serves as a representative of methods based accumulative attention scores. For all these methods, eviction ratio is set to 50%.
4.1.2 Results
The results in Table 1 demonstrate the effectiveness of our method in balancing accuracy and KV-Cache eviction. Our method consistently outperforms the baseline strategies in terms of accuracy across most tasks while achieving a higher mean eviction ratio. This indicates that our method is able to evict more tokens on average without significantly compromising accuracy.
In terms of accuracy, our method achieves competitive performance, closely matching the performance of vanilla Llama2-7B in most cases. For instance, on PIQA, COPA, and BoolQ, our method performs comparably to Llama2-7B, demonstrating minimal accuracy degradation. Compared to baselines, our method shows superior performance, particularly in tasks like BoolQ and PIQA, where both accuracy and eviction ratio surpass the baselines. This confirms the advantage of learnable mechanism for more efficient token retention strategies without sacrificing model performance.
Moreover, it is worth highlighting that our method achieves these results with minimal computational overhead, as mentioned in Section 4.1.1. The continual pre-training was conducted on only 5,000 samples and trained for just one epoch, demonstrating the lightweight nature of our approach. This efficiency can be attributed to the fact that our method does not need to learn new knowledge from scratch but rather focuses on learning effective token retention strategies, leveraging the existing capabilities of the pre-trained model.
4.2 Supervised Fine-tuning
Figure 3: This visualization highlights attention patterns in Llama2-7B after fine-tuning on the BoolQ dataset, using a selected sample to showcase multiple heads within both MHA and AG of different layers. In part (i), we visualize attention scores from several MHA heads across different layers before eviction. 1. MHA heads exhibit diverse attention patterns, especially in the first two layers, where attention is highly heterogeneous across heads. 2. While the first two layers show dense attention patterns, the subsequent layers become progressively sparser. 3. Bright-yellow vertical lines consistently appear at the same position across nearly all heads, especially beyond the first two layers, indicating critical tokens for inference. This aligns with the Heavy Hitters in H2O (Zhang et al., 2024), where a small portion of tokens significantly contributes to the attention scores. Our method’s effectiveness is demonstrated by the persistence of these bright lines in deeper layers, showing that crucial tokens are retained rather than evicted, ensuring their importance is preserved across the network. In part (ii), we visualize the attention scores from the attention-like structure within the AG mechanism. As the model layers deepen, the attention pattern shifts from high-resolution to lower-resolution, indicating that AG increasingly focuses on distilling in-context information. Thus, deeper AG layers no longer need such high resolution to capture global information, as the model has already refined and summarized it. This suggests potential optimizations, such as reducing the number of heads or dimensionality in deeper AG layers, to further improve efficiency.
4.2.1 Setup
Model and tasks. We utilize Llama2-7B (Touvron et al., 2023) as base model. To evaluate our approach, we select six widely recognized downstream tasks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), and OpenBookQA (Mihaylov et al., 2018). For each task, we fine-tune the model using the respective training set and evaluate its performance on the corresponding test set.
Baselines and Implementation Details. For the baseline, we apply LoRA (Hu et al., 2021) to fine-tune the models on each task, targeting the $(W_{q},W_{k},W_{v},W_{o})$ weights in the self-attention modules. In our method, in addition to applying LoRA to these parameters, the AG modules are also set to be learnable. As mentioned in Section 3.4, the following hyperparameters are used: $\tau=0.3$, $\gamma=0$, $\alpha=1$ or $0.5$, and $\beta=0.28$. We employ the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 5e-5, and train for two epochs on each dataset.
4.2.2 Results
As shown in Table 2, our method demonstrates a strong balance between accuracy and KV-Cache eviction across the six downstream tasks.
With $\alpha=1$, our method maintains competitive accuracy compared to the fine-tuned Llama2-7B baseline, with minimal drops (e.g., 82.15% vs. 82.92% on PIQA), while achieving a high mean eviction ratio of 60.00%. With $\alpha=0.5$, the eviction rate is reduced to 55.49%. But in some tasks like RTE and COPA, our method even surpasses the baseline accuracy. The average accuracy also exceeds the baseline (79.06% vs. 77.86%), suggesting that effectively evicting redundant tokens allows the model to focus on relevant information and improves performance.
Additionally, under the same hyperparameter settings, the performance varies across tasks. For instance, ARC-C is more challenging to evict compared to OpenBookQA, leading to a larger accuracy drop post-eviction. This highlights the importance of task-specific KV-Cache eviction policies.
4.3 Visualization
Table 3: Ablation study on various settings of AG, reporting accuracy (Acc.) and KV-Cache eviction ratio (%Eviction) under different configurations, with $\alpha=1$ in all settings. The results for (1) correspond to the setup described in Section 4.2, where the number of AG heads is 4, and the head dimension is 128. In (2-1) and (2-2), we explore the impact of the number of AG heads, with (2-1) using 2 heads and (2-2) using 1 head. The comparison between (1), (2-1), and (2-2) shows that reducing the number of heads leads to a drop in both accuracy and eviction ratio, indicating that the capacity of AG is closely tied to the number of heads. For (3-1) and (3-2), we assess the effect of reducing head dimensions for AG heads, where (3-1) has half the dimension size of (1) and (3-2) has 1/4. Comparing (1), (3-1), and (3-2) reveals that smaller dimensions reduce the eviction capability and accuracy, highlighting the importance of maintaining sufficient head dimensionality. Settings (4-1) and (4-2) examine using the hidden states and the AG module from the previous layer to inform the current layer’s eviction strategy. In (4-1), the first layer does not evict, and from the second layer onward, the eviction is determined by the previous layer. In (4-2), the first and second layers do not evict, following the observation from Section 4.3 that the first two layers are denser than subsequent layers. Both (4-1) and (4-2) show a slight decline in accuracy and eviction ratio compared to (1), but this approach introduces parallelism, offering a potential avenue for future optimizations. The setting (5) replaces the attention-like structure in AG with a simple linear layer to determine the eviction strategy. The comparison with (1) shows that linear layers almost cannot evict tokens effectively, reinforcing the necessity of leveraging global in-context information for successful eviction, as discussed in Section 3.3.
Metric PIQA ARC-C RTE COPA BoolQ OpenBookQA Avg.
(1) Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
%Eviction 66.16 48.31 65.47 45.40 67.46 67.17 60.00
(2-1) Acc. 81.88 57.63 65.70 91.00 87.52 77.40 76.86
%Eviction 63.92 36.38 62.73 24.38 65.22 63.57 52.70
(2-2) Acc. 82.15 53.90 62.45 89.00 87.31 77.40 75.37
%Eviction 58.97 31.47 59.77 20.32 63.02 59.17 48.79
(3-1) Acc. 81.45 53.36 58.84 88.00 86.73 78.40 74.46
%Eviction 61.75 33.55 61.34 19.24 64.59 59.59 50.01
(3-2) Acc. 83.03 53.90 59.93 89.00 87.16 76.40 74.90
%Eviction 58.68 24.23 32.23 12.28 59.40 55.54 40.39
(4-1) Acc. 81.66 55.25 66.06 88.00 86.85 78.00 75.97
%Eviction 49.52 36.92 46.85 28.74 56.02 60.32 46.40
(4-2) Acc. 82.75 55.93 79.06 82.00 86.33 78.40 77.41
%Eviction 53.31 44.38 51.20 47.95 61.98 61.73 53.43
(5) Acc. 82.54 54.58 57.40 81.00 87.71 74.80 73.01
%Eviction 1.06 0.46 0.81 0.26 1.38 1.16 0.86
In this section, we present visualizations to highlight key characteristics of our AG mechanism. After fine-tuning on specific tasks, we visualize the model’s MHA and AG weights for selected samples, as shown in Figure 3. Additional visualizations can be found in Section A.2.
4.4 Ablation
In this ablation study, we explore the effects of various configurations on the AG mechanism, focusing on the number of AG heads, AG head dimensions, and eviction strategies. For details of each setting and the corresponding performance, refer to Table 3. Key findings include the impact of reducing the number of heads, as shown in (2-1) and (2-2), and the reduction of head dimensions in (3-1) and (3-2), both of which result in lower accuracy and eviction ratios. In (4-1) and (4-2), we examine eviction strategies where the current layer’s eviction is based on the previous layer’s hidden states and AG module, which introduces parallelism but slightly reduces performance. Finally, replacing the AG with a linear layer in (5) demonstrates the necessity of using an attention-like structure for effective token eviction. For more ablation, please refer to Table 5.
5 Conclusion
In conclusion, the proposed Attention-Gate mechanism offers a flexible and adaptive solution to KV-Cache eviction in large language models. By dynamically identifying and discarding less important tokens, Attention-Gate addresses the limitations of static and attention-score-based strategies, providing efficient context-aware eviction. This mechanism integrates seamlessly with pre-trained models and can be easily tuned, making it a practical and effective method for enhancing both performance and memory efficiency in various tasks.
Appendix A Appendix
A.1 Results of Continual Pre-training on Mistral
We conducted continual pre-training on Mistral using 5,000 samples from RedPajama (Computer, 2023), and the results are shown in Table 4. Compared to the performance of Llama2-7B presented in Table 1, Mistral’s performance slightly declined. We hypothesize that this may be due to the distribution of RedPajama’s data being less suited to Mistral. Additionally, this raises the question of whether KV-Cache eviction is model-dependent, and whether its effectiveness is related to the model’s expressive power. Although the parameter counts of Mistral-7B and Llama2-7B are similar, Mistral-7B significantly outperforms Llama2-7B. This could suggest that Mistral is utilizing more tokens or scoring them with finer granularity, which results in fewer redundant tokens and thus makes eviction less effective.
Table 4: Performance comparison between Mistral-7B and Ours across various tasks.
Metric PIQA ARC-C ARC-E RTE COPA BoolQ HellaSwag Avg.
Mistral-7B Acc. 80.09 42.37 63.14 48.01 76 64.22 73.02 63.84
Ours Acc. 75.9 34.24 55.2 48.01 65 62.2 67.91 58.35
Eviction 37.14 39.48 37.80 40.93 45.27 44.68 50.92 42.32
A.2 More Visualization
Figure 4: The complete version of Figure 3.
Figure 5: More samples for the same scenario of Figure 3.
Figure 4 provides a comprehensive view of the layers and attention heads from Figure 3. Additionally, four sample scenarios for the same setup are presented in Figure 5.
A.3 More Ablation
Table 5: Exploring the impact of the number of recent tokens, viewed from the perspective of the attention matrix and considering slanted retention patterns. (1) corresponds to the setup described in Section 4.2, where only the current token is retained, and thus reflecting only the diagonal retention in the attention matrix. For (6-1) and (6-2), the number of recent tokens retained is set to 5 and 10, respectively. The results suggest that increasing the number of recent tokens does not necessarily enhance performance under the AG framework. Further exploration of how to manage recent tokens, such as applying learnable weighted strategies, could be an interesting direction for future work.
Metric PIQA ARC-C RTE COPA BoolQ OpenBookQA Avg.
(1) Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
%Eviction 66.16 48.31 65.47 45.40 67.46 67.17 60.00
(6-1) Acc. 83.08 50.85 65.34 82.00 87.31 73.20 73.63
%Eviction 65.15 40.96 64.29 21.37 67.49 63.69 53.83
(6-2) Acc. 81.61 53.56 60.29 82.00 87.37 74.20 73.17
%Eviction 65.66 44.48 65.14 24.28 68.18 63.44 55.20
In this section, we present additional ablation results in Table 5.

### v3

@@
 The KV-Cache technique has become the standard for the inference of large language models (LLMs).
-It caches states of self-attention to avoid recomputation.
-Yet, it is widely criticized that KV-Cache can become a bottleneck of the LLM inference system, especially when confronted with ultra-large models and long-context queries.
-A natural remedy is to discard the KV-Cache for less important tokens, with StreamingLLM (Xiao et al., 2024) as an example, but the used static eviction strategies cannot flexibly adapt to varying contexts.
-Remedies like H2O (Zhang et al., 2024) leverage accumulative attention scores to perform dynamic eviction but suffer from the attention bias issue in capturing contextual information.
-This paper bridges this gap by devising a parameterized KV-Cache eviction mechanism, dubbed as Attention-Gate, which accepts the whole context as input and yields eviction flags for each token to realize in-context eviction.
-The subsequent self-attention module proceeds according to the flags and only the KV states for the remaining tokens need to be cached.
-The Attention-Gates can vary among different heads and layers and be trivially plugged into pre-trained LLMs, tuned by cost-effective continual pre-training or supervised fine-tuning objectives to acquire what to discard.
-The computational and memory overhead introduced by Attention-Gates is minimal.
-Our method is validated across multiple tasks, demonstrating both efficiency and adaptability.
-After a highly efficient continual pre-training, it achieves higher average accuracy and evicts more tokens compared to traditional training-free methods.
-In supervised fine-tuning, it not only evicts many tokens but also outperforms LoRA-finetuned LLMs on some datasets, such as RTE, where it improves accuracy by 13.9% while evicting 62.8% of tokens, showing that effective eviction of redundant tokens can even enhance performance.
+Yet, it is widely criticized that KV-Cache can become a bottleneck of the LLM inference system.
+This paper enables a novel dynamic KV-Cache eviction policy by injecting a lightweight module called Attention-Gate to the model.
+It accepts the global context as input and yields eviction flags for each token.
+The self-attention modules in the model proceed according to the flags and cache only a subset of the KV states for next token prediction.
+The Attention-Gates can yield various flags for different heads and layers and be easily tuned on top of a pre-trained LLM via continual pre-training or supervised fine-tuning.
+The computational and memory overhead introduced by Attention-Gates can be minimal.
+We empirically evaluate the proposed approach across multiple scenarios, showing that effective eviction of redundant tokens can not only improve efficiency but also enhance performance.
+Machine Learning, ICML
 1 Introduction
-Large language models (LLMs) (Dubey et al., 2024; Team et al., 2024; Chiang et al., 2023) have achieved remarkable success across a wide range of natural language processing tasks.
+Large language models (LLMs) (Dubey et al., 2024; Team et al., 2024; Chiang et al., 2023) have achieved remarkable success across a wide range of tasks.
 A key technique that has enabled efficient LLM inference is KV-Cache, which stores transient attention keys and values to avoid recomputation.
@@
 A common trend observed is that deeper layers tend to mask more KV-Cache states, with some in deeper layers being entirely masked.
-To overcome these challenges, we introduce a parameterized KV-Cache eviction mechanism named the Attention-Gate (AG), to perform reliable in-context eviction.
-AG is positioned before a self-attention layer within the model.
-It takes a sequence of token features as input and generates eviction flags for the tokens, indicating whether a token should be excluded from subsequent self-attention computations.
+To overcome these challenges, we introduce a parameterized KV-Cache eviction mechanism named Attention-Gate (AG), to perform reliable in-context eviction.
+AGs are positioned before self-attention layers within the model.
+They take a sequence of token features as input and generate eviction flags for the tokens, indicating whether the token should be excluded from subsequent self-attention computations.
 Tokens that are evicted do not require their KV states to be cached.
@@
 AG is empirically shown to enjoy high training efficiency, e.g., only four NVIDIA 4090 GPUs and a dataset of 5,000 samples are required for continual pre-training when applying AGs to LLaMA2-7B (Touvron et al., 2023).
-This alleviates concerns about the computational overhead related to trainable eviction strategies (Zhang et al., 2024; Chen et al., 2024) and amplifies the performance superiority of our approach over existing training-free approaches.
+This alleviates concerns about the computational overhead related to trainable eviction strategies (Zhang et al., 2024; Chen et al., 2024) and amplifies the merits of our approach over existing training-free approaches.
 As illustrated in Figure 1, AG generates different eviction strategies across different layers and attention-heads for different tokens, demonstrating its adaptability to the diverse requirements of each component in the model.
@@
 In supervised fine-tuning (SFT), our method not only evicts a significant number of redundant tokens but also maintains or surpasses the performance of LoRA-finetuned LLMs.
-For example, on the RTE dataset, our approach improves accuracy by 13.9% while evicting 62.8% of tokens, demonstrating that selective token eviction can enhance performance.
+For example, on the RTE dataset (Bar-Haim et al., 2006), our approach improves accuracy by 13.9% while evicting 62.8% of tokens, demonstrating that selective token eviction can enhance performance.
 In summary, the Attention-Gate mechanism provides a scalable and efficient solution for KV-Cache management, addressing the limitations of traditional training-free methods.
-Figure 2: An overview of Attention-Gate (AG) for KV-Cache eviction.
-AG is a lightweight learnable module placed before each MHA layer.
-Given the input hidden states, it determines for each head whether to retain or discard the key and value tokens in the KV-Cache.
-In the attention weights, this corresponds to masking out columns for the evicted keys, while keeping the diagonal intact to ensure the query interacts with its own key.
 2 Related Work
@@
 A significant bottleneck lies in the attention mechanism, which demands considerable computational and memory resources, especially for long sequences.
-Computation and memory challenges.
-The computational cost of attention is driven by its quadratic complexity, while the memory burden stems from the need to store the KV-Cache during inference to avoid recomputation.
-Although several works have focused on reducing the computational load of attention mechanisms, many fail to address the memory constraints that remain a critical issue.
-Techniques like Reformer (Kitaev et al., 2020) and FlashAttention (Dao et al., 2022; Dao, 2023) mitigate memory usage for long sequences but still require substantial cache storage.
-On the other hand, approaches such as MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023) prioritize memory compression without tackling computational complexity.
-These solutions tend to focus on either memory or computational efficiency, rarely addressing both simultaneously.
-KV-Cache eviction strategies.
+Traditional KV-Cache eviction strategies.
 To address both memory and computational challenges, KV-Cache eviction has emerged as an effective strategy.
-Existing approaches can be categorized into static and adaptive methods based on accumulative attention scores.
+Existing approaches predominantly rely on parameter-free heuristics.
 Static strategies, such as those used in Sparse Transformers (Child et al., 2019), employ fixed pruning patterns, such as Strided and Fixed Attention.
@@
 However, none of these approaches fully resolve the underlying problem.
+Despite these limitations, some studies (Adams et al., 2024) suggest that H2O remains an optimal solution in many scenarios.
 More adaptive strategies.
@@
 3.1 Preliminary
-Multi-Head Attention (MHA) (Vaswani et al., 2017) is a core component of the Transformer architecture, widely used in most LLMs.
-MHA enables the model to capture dependencies across different tokens in a sequence by computing self-attention across multiple attention-heads.
-Each head attends to different parts of the input sequence independently, allowing the model to capture various aspects of the relationships between tokens.
-The outputs of these attention heads are then concatenated and projected through a final output matrix.
-KV-Cache is employed to store the key and value representations of tokens from previous time steps during inference.
-This prevents redundant recomputation of these representations for every new token generated, significantly speeding up auto-regressive generation.
-The inference process in auto-regressive transformers, such as Llama (Touvron et al., 2023), can be divided into two stages: prefilling and decoding.
-1.
-In the prefilling stage, the model processes the entire input sequence, generating key-value pairs for each attention-head in each layer.
-These key-value pairs are stored in the KV-Cache for future reuse.
-2.
-In the decoding stage, when generating the next token, the model only computes the key and value for that new token and appends them to the cached key-value pairs.
-This reduces the computational load, as the previously cached key-value pairs do not need to be recomputed.
-The process continues token by token until the sequence generation is complete.
-KV-Cache plays a critical role in improving the efficiency of LLM inference, especially in scenarios where long sequences are processed.
+Multi-Head Attention (MHA) (Vaswani et al., 2017) is a core component of the Transformer architecture, as used by most LLMs.
+MHA enables the model to capture dependencies across different tokens in a sequence.
+Specifically, for an input sequence $X\in\mathbb{R}^{n\times d}$, where $n$ represents the sequence length and $d$ denotes the dimensionality of the hidden states, the output of MHA is computed as:
+$\text{MHA}(X)=\left[\text{H}_{1}(X),\text{H}_{2}(X),\dots,\text{H}_{h}(X)\right]W^{O}\ ,$ (1)
+where $\{\text{H}_{i}\}_{i=1}^{h}$ refers to the $h$ attention heads and
+$\text{H}_{i}(X)$ $=\text{Attn}\left(XW^{Q}_{i},XW^{K}_{i},XW^{V}_{i}\right)$
+$=\text{Attn}\left(Q_{i},K_{i},V_{i}\right)$ (2)
+$=\text{Softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{k}}}-\text{INF}(\mathbbm{1}-M_{i})\right)V_{i}=A_{i}V_{i}.$
+Here, $W^{Q}_{i},W^{K}_{i}\in\mathbb{R}^{d\times d_{k}}$, $W^{V}_{i}\in\mathbb{R}^{d\times d_{v}}$, and $W^{O}\in\mathbb{R}^{hd_{v}\times d}$ are learned projection matrices.
+INF is a large constant, $\mathbbm{1}$ is a matrix of ones, $M_{i}$ is the mask applied to head Hi, and $A_{i}$ represents the attention scores for head Hi.
+KV-Cache is employed during the inference of auto-regressive transformers, which stores the key and value information from previous time steps, allowing efficient reuse and reducing recomputation.
+The inference process can be divided into two stages: prefilling and decoding.
+In the prefilling stage, the input sequence $X^{(\leq n)}=\left[x^{(1)},x^{(2)},\dots,x^{(n)}\right]\in\mathbb{R}^{n\times d}$ passes through MHA, and the corresponding key-value pairs $K^{(\leq n)}_{i}$ and $V^{(\leq n)}_{i}$ for head Hi are stored in KV-Cache.
+These are expressed as:
+$K_{i}^{(\leq n)}=\left[k_{i}^{(1)},\cdots,k_{i}^{(n)}\right],\ V_{i}^{(\leq n)}=\left[v_{i}^{(1)},\cdots,v_{i}^{(n)}\right]\ ,$
+where $k_{i}^{(t)}=x^{(t)}W_{i}^{K}$ and $v_{i}^{(t)}=x^{(t)}W_{i}^{V}$. After prefilling, the next token $x^{(n+1)}$ is generated.
+In the decoding stage, $x^{(n+1)}$ is input to generate $x^{(n+2)}$ for the first step.
+During this process, only $k_{i}^{(n+1)},v_{i}^{(n+1)}$ need to be computed for head Hi.
+These are then concatenated with the cached $K_{i}^{(\leq n)}$ and $V_{i}^{(\leq n)}$ to form $K_{i}^{(\leq n+1)}$ and $V_{i}^{(\leq n+1)}$, which are used to complete the current MHA computation and update the KV-Cache.
+The process repeats token by token until the sequence generation is complete.
+KV-Cache plays a critical role in improving the efficiency of LLM inference.
 However, the size of the KV-Cache grows with the input sequence length, leading to substantial memory overhead.
-Efficiently managing this cache while maintaining model performance has become a key challenge in scaling LLMs to longer contexts.
+Efficiently managing KV-Cache while maintaining model performance has become a key challenge in scaling LLMs to longer contexts.
 3.2 Limitations of Traditional Eviction Strategies
 Lack of flexibility.
-Flexibility in KV-Cache eviction strategies spans across various dimensions, including token-specific, attention-head-specific, layer-specific, task-specific, and model-specific adaptability.
-Static eviction policies offer no such flexibility, requiring manual intervention to adjust them.
-While methods based on accumulative attention scores (e.g., H2O (Zhang et al., 2024)) have improved upon static strategies by introducing token-level and head-level adaptability, they are still limited.
-Specifically, in H2O, the eviction ratio remains uniform across all attention heads, restricting head-level flexibility.
-Flexibility is crucial in eviction strategies because models process varying types of data across different contexts, requiring the ability to selectively retain important tokens while discarding others.
-Without this adaptability, the model may inefficiently store redundant information, increasing memory usage and slowing down inference times.
+Static KV-Cache eviction strategies, such as those used in StreamingLLM (Xiao et al., 2024), lack adaptability across key dimensions, including token-specific, attention-head-specific, layer-specific, task-specific, and model-specific contexts.
+H2O (Zhang et al., 2024) addresses some of these limitations by introducing token-level and head-level adaptability.
+However, it still applies a uniform eviction ratio across all attention heads, which limits its granularity at the head level.
+This isotropic approach overlooks the significant variation in attention scores across different heads, as demonstrated by FastGen (Ge et al., 2023).
+This observation underscores the need for eviction strategies that are adaptive across multiple dimensions to better handle the diverse contexts processed by the model.
+Without finer-grained flexibility, existing training-free strategies risk retaining unnecessary information, leading to higher memory usage and reduced efficiency.
 Absence of global statistics.
-Eviction strategies that rely on accumulative attention scores, such as H2O, base their calculations on local statistics, which often results in attention bias issues (Oren et al., 2024).
-This local approach may misjudge a token’s importance, especially when the context spans long sequences.
-While several methods (Chen et al., 2024; Jo & Shin, 2024) have proposed solutions to mitigate this bias, they are still fundamentally rooted in local statistics, addressing the issue only partially.
-A more effective solution would incorporate global statistics to ensure that the eviction process is based on a more comprehensive understanding of the context, thereby reducing biases and improving the accuracy of token retention.
+KV-Cache eviction strategies should ideally be context-aware, as the importance of the same token can vary significantly depending on the surrounding context.
+To accurately discard redundant or less important tokens, it is necessary to consider the global statistics of the context.
+Strategies relying on accumulative attention scores, such as H2O, NACL, and A2SF (Chen et al., 2024; Jo & Shin, 2024; Zhang et al., 2024), are primarily based on local statistics of the context.
+While methods like NACL and A2SF attempt to mitigate the limitations of local context by introducing various adjustments to reduce the misjudgment of token importance and the resulting biased retention (Oren et al., 2024), they fail to address the root issue.
+The core problem lies in the lack of consideration of the global context when determining which tokens to evict.
 Inefficiency.
-Efficiency is another challenge for traditional methods such as H2O and FastGen (Ge et al., 2023).
-These methods operates in a token-by-token manner for each decoding step, sequentially deciding which key-value pairs to evict from the cache.
-Additionally, H2O determines which tokens to evict after computing the attention scores, meaning some computation is wasted on tokens that are later discarded.
-This sequential approach hinders efficiency, especially when handling large batches of data or when processing long sequences.
-Our Attention-Gate mechanism is designed to address the aforementioned limitations.
+Methods like H2O and FastGen (Ge et al., 2023) are inefficient due to their sequential, token-by-token eviction processes at each decoding step.
+H2O, for example, computes attention scores before deciding which tokens to evict, wasting computation on soon-to-be discarded tokens.
+Our method tries to address the aforementioned limitations.
+Figure 2: An overview of Attention-Gate (AG) for KV-Cache eviction.
+AG is a lightweight learnable module placed before each MHA layer.
+Given the input hidden states, it determines for each head whether to retain or discard the key and value tokens in the KV-Cache.
+In the attention weights, this corresponds to masking out columns for the evicted keys, while keeping the diagonal intact to ensure the query interacts with its own key.
 3.3 Attention-Gate
-The Attention-Gate (AG) is a lightweight, trainable module designed to determine which tokens in the KV-Cache of each attention-head should be retained or discardedAG is positioned before the Multi-Head Attention (MHA) layer and operates by generating eviction flags for each token, guiding both the computation of attention scores and the management of the KV-Cache.
-Input and output.
+The Attention-Gate (AG) is a lightweight, trainable module positioned before the MHA layer, generating eviction flags for the tokens.
+The flags determine which tokens in the KV-Cache of each head of the MHA should join the computation of attention scores and be retained in the KV-Cache.
+Main architecture.
+AG consists of (i) an attention-like structure and (ii) a gating mechanism.
+(i) A modified version of the standard MHA in Equation 1, denoted as $\text{MHA}^{\prime}$, is used to facilitate global information exchange across all tokens in the sequence.
+To distinguish it from the vanilla MHA, all symbols in $\text{MHA}^{\prime}$ are marked with a prime ($\prime$).
+Notably, $\text{MHA}^{\prime}$ has much fewer heads compared to MHA, i.e., $h^{\prime}<h$, and its projection matrix $W^{O^{\prime}}\in\mathbb{R}^{h^{\prime}d^{\prime}_{k}\times h}$ differs slightly from $W^{O}\in\mathbb{R}^{hd_{v}\times d}$.
+(ii) The gating mechanism, denoted as G, is then introduced to enable adaptive eviction policies for attention heads within the MHA layer.
+Input & output.
 AG takes the hidden states as input and generates binary flags as output, one for each token in every attention head.
-These flags indicate whether the corresponding token’s key-value pairs should be stored or discarded in the KV-Cache.
-Specifically, the input is a hidden state matrix $X\in\mathbb{R}^{n\times d}$, where $n$ represents the sequence length, and $d$ denotes the dimensionality of the hidden states.
-The output consists of independent binary decisions for the retention or eviction of each token’s key-value pairs across all attention heads in the MHA module.
-Local or global?
-The structure of AG offers two choices: using either local or global information to guide the eviction policy.
-1.
-Local information: The simplest approach leverages a linear layer where each token only considers its own hidden state without taking other tokens in the sequence into account.
-This method is computationally efficient and straightforward, making it appealing for its simplicity.
-2.
-Global information: On the other hand, a more comprehensive approach adopts an attention-like structure, where tokens can aggregate information from across the entire sequence.
-This enables the eviction policy to reflect the global in-context information, making decisions that are informed by the broader context.
-Naturally, this method is computationally heavier than the local approach, but it offers a more accurate reflection of the sequence’s token importance.
-In our design, we opted for the attention-like structure because global context is essential for making effective eviction decisions.
-As demonstrated in Section 4.4, using an attention-like structure significantly outperforms the linear layer, as the latter struggles to properly evict tokens.
-This outcome is expected, as determining whether a token is redundant or crucial for inference often requires a global understanding of the sequence.
-A local approach, like a linear layer focusing on individual tokens, lacks the broader perspective necessary to make accurate decisions about which tokens are important.
-This finding also underscores the limitations of methods like H2O, which rely on local statistics rather than global ones, as mentioned in Section 3.2.
-Softmax or Sigmoid?
-After gathering information through the attention-like structure, the next step is to compute the eviction probability for each token.
-Common activation functions such as Softmax and Sigmoid can be used to generate these probabilities.
-In our case, we choose the Sigmoid function for a key reason: while Softmax normalizes probabilities across the entire sequence, introducing competition between tokens, Sigmoid treats each token independently.
-This independence is crucial because we do not want a token’s likelihood of being evicted to depend on the eviction probabilities of other tokens in the sequence.
-Eviction flags.
-After computing the eviction probabilities for each token, we apply a threshold $\tau$.
-If the probability of retaining a token exceeds $\tau$, that token is retained in the KV-Cache; otherwise, it is discarded.
+Specifically, given an input sequence $X\in\mathbb{R}^{n\times d}$, the output of AG is computed as:
+$\text{AG}(X)=\text{G}\left(\text{MHA}^{\prime}(X),\tau\right)\ \in\{0,1\}^{n\times h},$ (3)
+where $\tau$ is a hyperparameter. The gate G outputs 1 or 0 based on the score $s$ of a token obtained from MHA:
+$\text{G}(s,\tau)=\begin{cases}1,&\text{if }\text{Sigmoid}(s)>\tau\\ 0,&\text{otherwise}\end{cases}\ .$ (4)
+The values $k_{i}^{(t)}$ and $v_{i}^{(t)}$ for token $x^{(t)}$ and attention-head Hi are retained if $\text{AG}(X)_{i}^{(t)}=1$; otherwise, they are discarded.
+Based on this, the mask $M_{i}=\big{[}m_{i}^{(j,t)}\big{]}$ for the attention scores $A_{i}$ of head Hi in Section 3.1 is defined as:
+$m_{i}^{(j,t)}=\begin{cases}0,&\text{if }j<t\\ 1,&\text{if }j=t\\ \text{AG}(X)_{i}^{(t)}\ \ ,&\text{otherwise}\end{cases}\ .$ (5)
 In the attention matrix, the columns corresponding to evicted tokens are masked out.
 However, the diagonal elements, where a token attends to itself, are always preserved.
-This guarantees that each token continues to interact with its own key.
+This ensures that each token maintains interaction with its own key, guaranteeing self-attention remains intact.
+In this way, the AG module selectively determines which tokens are retained or discarded for each attention head, based on the global information captured by $\text{MHA}^{\prime}$ and the gating mechanism.
+3.4 Design Rationale
+Local or global?
+Our AG design, introduced in Section 3.3, utilizes a multi-head attention mechanism (i.e.
+MHA) that inherently incorporates global information.
+This approach allows the model to aggregate context across the entire sequence, enabling eviction decisions to accurately reflect the broader in-context information.
+Such a design is particularly effective for assessing the importance of tokens, as determining redundancy or relevance often requires a global understanding of the sequence.
+Alternatively, a simpler approach can be used, such as employing a linear transformation to generate eviction flags.
+This method relies solely on local information, where each token only considers its own hidden state without incorporating information from other tokens in the sequence.
+While the local approach has the advantage of being computationally cheaper, as shown in Table 3, its performance and efficiency are not guaranteed.
+This limitation highlights the challenges faced by methods rely on local statistics of the context, as discussed in Section 3.2.
+Softmax or Sigmoid?
+In our design, as described in Equation 4, we use the Sigmoid function to calculate the eviction probability for each token.
+However, it is worth noting that Softmax is another common choice for generating probabilities, as seen in applications like the gating mechanisms of Mixture of Experts in LLMs.
+While both functions are viable, they offer different trade-offs.
+The Sigmoid function, as used in our design, computes probabilities independently for each token.
+This independence is important because it ensures that a token’s eviction likelihood is not directly influenced by the probabilities of other tokens in the sequence, allowing for a more flexible and precise eviction policy.
+On the other hand, Softmax normalizes the probabilities across the entire sequence, which introduces competition between tokens.
+This can be beneficial in scenarios where the relative importance of tokens needs to be considered.
+However, in the context of eviction, where we want to assess each token individually without direct dependencies on others, Softmax might not be the ideal choice.
 Prefilling or decoding?
@@
 Using AG in decoding could slow down inference and increase the cache memory footprint, as additional keys and values from the attention-like structure would need to be stored.
-In summary, AG dynamically determines which tokens should be retained, optimizing both KV-Cache usage and attention score computation by utilizing an attention-like structure, Sigmoid activation function, and a threshold-based decision for token eviction.
 3.5 Training Implementation
-To train the Attention-Gate (AG) effectively, this section outlines the key components of the training process. For more comprehensive details, please refer to Section 4 and Appendix A.
+To train the Attention-Gate (AG) effectively, this section outlines the key components of the training process. For more details, please refer to Section 4 and Appendix A.
 Eviction Loss.
 To encourage the eviction of unnecessary tokens, we introduce a dedicated loss function called the Eviction Loss.
-This loss encourages the model to discard as many tokens as possible, which is defined as:
-$\ell_{\text{evict}}=\alpha\cdot\frac{1}{L}\sum_{l=1}^{L}\text{ReLU}\left(\overline{\text{AG}}_{l}-\beta\right),$ (1)
-where $L$ is the number of layers, and $\overline{\text{AG}}_{l}$ represents the average output of the AG module in layer $l$.
-In this formula, $\alpha$ adjusts the intensity of KV-Cache eviction, while $\beta\in[0,\tau]$ ensures that eviction does not become overly aggressive.This loss function works alongside the auto-regressive loss to balance token eviction with maintaining model performance.
-The loss allows for adaptive eviction across layers and attention-heads.
-As shown in Figure 1, after training, the model learns to increase token eviction in deeper layers, which is not achievable with training-free methods like H2O (Zhang et al., 2024) or StreamingLLM (Xiao et al., 2024).
+The loss function encourages the model to maintain the eviction ratio close to a target value $\beta$, which is defined as:
+$\ell_{\text{evict}}=\alpha\cdot\left|\overline{\text{AG}}-\beta\right|,$ (6)
+where $\overline{\text{AG}}$ represents the average output of all AG modules.
+In this formula, $\alpha$ adjusts the intensity of KV-Cache eviction, while $\beta\in[0,\tau]$ ensures that eviction does not become overly aggressive.
+Eviction Loss allows for adaptive eviction across layers and attention-heads.
+This loss function works alongside the auto-regressive loss to balance token eviction with maintaining model performance.
 Initialization.
 We initialize the AG parameters using Xavier initialization (Glorot & Bengio, 2010) to provide a stable starting point for learning.
-Additionally, a small constant $\gamma\geq 0$ is added inside Sigmoid, ensuring that the initial retention probabilities are close to 1.
-This encourages the model to retain most tokens early in training, allowing it to learn which tokens are important more gradually.
+Additionally, a small constant $\gamma\geq 0$ can optionally be added inside Sigmoid in Equation 4, ensuring that the initial retention probabilities are close to 1.
+This encourages the model to retain most tokens early in training.
 Handling non-differentiability.
@@
 This approach ensures smooth gradients and enables effective training of the AG while preserving its binary behavior during the forward pass.
+Table 1: Performance comparison of Llama2-7B and various KV-Cache eviction strategies after continual pre-training.
+For baselines, ($W_{q},W_{k},W_{v},W_{o}$) are made trainable, while in our method, the AG module is also trainable.
+Higher values indicate better performance for all metrics.
+Acc.
+refers to accuracy.
+%Evict.
+refers to the mean KV-Cache eviction ratio, representing the percentage of tokens evicted from KV-Cache.
+The eviction ratio is fixed at 50% for the baseline methods.
+In contrast, our method achieves better performance (average accuracy and score) while maintaining a higher average %Evict..
+Metric PIQA ARC-C ARC-E RTE COPA BoolQ HellaSwag MMLU Avg. Metric LongBench
+Llama2-7B-cpt Acc. 72.69 32.88 50.62 50.54 57.00 64.77 42.19 26.64 49.67 Score 23.42
+StreamingLLM Acc. 72.42 31.53 49.74 50.90 54.00 61.31 37.75 26.66 48.04 Score 4.61
+H2O Acc. 72.20 30.85 49.38 51.99 55.00 62.42 41.45 26.45 48.72 Score 4.85
+Ours Acc. 76.33 32.20 48.32 50.18 59.00 60.46 64.23 28.54 52.41 Score 13.71
+%Evict. 43.12 46.54 45.15 48.60 55.37 50.16 61.10 70.36 52.55 %Evict. 68.55
+3.6 Complexity Analysis
+AG module. For input $X\in\mathbb{R}^{n\times d}$, the FLOPs for AG are:
+$\operatornamewithlimits{\text{FLOPs}}_{\text{AG}}=\mathcal{O}(n^{2}d_{k}^{\prime}h^{\prime})\ .$
+MHA module. Without AG, original MHA’s FLOPs are:
+$\operatornamewithlimits{\text{FLOPs}}_{\text{original MHA}}=\mathcal{O}(n^{2}d_{k}h)\ .$ (7)
+After AG processing, $t\%$ of the KV-Cache tokens are discarded, leaving $(1-t\%)$ for attention computation:
+$\operatornamewithlimits{\text{FLOPs}}_{\text{MHA after AG}}=\mathcal{O}\big{(}(1-t\%)n^{2}d_{k}h\big{)}\ .$
+Combined AG & MHA, the total FLOPs are:
+$\operatornamewithlimits{\text{FLOPs}}_{\text{AG \& MHA}}=\mathcal{O}\big{(}n^{2}\max\big{(}d_{k}^{\prime}h^{\prime},(1-t\%)d_{k}h\big{)}\big{)}\ .$ (8)
+Efficiency.
+The reduction in FLOPs depends on three factors: (i) Reduction in token count ($t\%$): Higher values of $t\%$ result in a larger reduction in the quadratic term of the original MHA.
+(ii) Head configuration ($h^{\prime}<h$): The AG module must have significantly fewer heads ($h^{\prime}$) compared to the original MHA ($h$) to ensure its overhead is small.
+(iii) Head dimension ratio ($d_{k}^{\prime}<d_{k}$): A smaller head dimension ($d_{k}^{\prime}$) for AG further reduces its contribution to total FLOPs.
 4 Experiments
 This section consists of three main parts.
-First, we evaluate the performance of AG in two scenarios: continual pre-training (CPT) and supervised fine-tuning (SFT).
-For CPT, we compare AG with classic methods on performance and eviction rates (Table 2).
-For SFT, we benchmark AG against vanilla LoRA fine-tuning to highlight its task-specific adaptability (Section 4.2).
+First, we evaluate the performance of AG in two scenarios: continual pre-training (CPT) and supervised fine-tuning (SFT) (Section 4.1 & 4.2).
 Second, we provide visualization of selected examples to demonstrate the core characteristics of AG (Section 4.3).
 Finally, we conduct ablation studies to provide further insights into the effectiveness of AG (Section 4.4).
+Additional results are provided in Section A.1.
 4.1 Continual Pre-training
+4.1.1 Setup
+Models & datasets.
+We use Llama2-7B (Touvron et al., 2023) as our primary base model.
+Additionally, we validate the feasibility of our approach on Mistral-7B (Jiang et al., 2023), with results provided in Table 5.
+For continual pre-training, we select a subset of the RedPajama dataset (Computer, 2023), comprising approximately 5,000 samples  [^1] Specifically, we sampled 4,997 samples proportionally from each subset of the RedPajama dataset., to serve as the training set.
+To assess the effectiveness of our method, we evaluate it on widely recognized benchmarks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021) and LongBench (Bai et al., 2023).
+All evaluations are conducted in a zero-shot setting, with performance assessed using OpenCompass (Contributors, 2023).
+Training details & baselines.
+For the Llama2-7B model, the following hyperparameters from Section 3.5 were used: $\tau=0.5$, $\gamma=2$, $\alpha=5$, and $\beta=0.4$.
+The model was trained for a single epoch.
+The evaluation metrics include performance (accuracy and score) and the mean eviction ratio for all KV-Cache.
+For KV-Cache eviction strategies, we use StreamingLLM (Xiao et al., 2024) as a representative of static strategies, while H2O (Zhang et al., 2024) represents methods based on accumulative attention scores  [^2] Adams et al.
+(2024) suggest that H2O is an optimal solution in many scenarios.
+Due to its strong performance and the difficulty of reproducing other training-free methods, no additional training-free baselines were included..
+For baseline methods, eviction ratio is fixed at 50%, and ($W_{q},W_{k},W_{v},W_{o}$) are made trainable using LoRA (Hu et al., 2021).
+In our method, AG is also trainable  [^3] We tested a version where only AG is trainable, with other parameters frozen.
+Results are shown in Table 4..
+4.1.2 Results
+The results in Table 1 demonstrate the effectiveness of our method in balancing performance and KV-Cache eviction.
+Our method consistently outperforms the baseline strategies in terms of performance across most tasks while achieving a higher mean eviction ratio.
+Moreover, it is worth highlighting that our method achieves these results with minimal computational overhead, as detailed in Section 4.1.1.
+The continual pre-training was conducted on only 5,000 samples and trained for just one epoch, demonstrating the lightweight nature of our approach.
+This efficiency can be attributed to the fact that our method does not need to learn new knowledge from scratch but rather focuses on learning effective token retention strategies, leveraging the existing capabilities of the pre-trained model.
+Table 2: Performance of Llama2-7B with LoRA fine-tuning and our method on six downstream tasks.
+In addition to the LoRA fine-tuning targets, our method makes the AG modules learnable.
+Two settings for $\alpha$ (0.5 and 1) are tested.
+Our method maintains comparable or better accuracy while achieving a higher eviction ratio, demonstrating its task-specific adaptability in managing token eviction.
+Metric PIQA ARC-C RTE COPA BoolQ OBQA Avg.
+Fine-tuned Acc. 82.92 60.34 64.98 92.00 88.10 78.80 77.86
+Llama2-7B
+Ours Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
+$(\alpha=1)$Evict. 66.16 48.31 65.47 45.40 67.46 67.17 60.00
+Ours Acc. 81.50 57.63 74.01 95.00 87.00 79.20 79.06
+$(\alpha=0.5)$Evict. 64.96 36.45 62.80 34.77 67.31 66.65 55.49
+4.2 Supervised Fine-tuning
+4.2.1 Setup
+Model & tasks.
+We utilize Llama2-7B (Touvron et al., 2023) as base model.
+To evaluate our approach, we select six widely recognized downstream tasks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), and OpenBookQA (Mihaylov et al., 2018).
+For each task, we fine-tune model using the respective training set and evaluate its performance on the corresponding test set.
+Implementation details.
+The selection of trainable parameters follows Section 4.1.1.
+The hyperparameters are $\tau=0.3$, $\gamma=0$, $\alpha=1$ or $0.5$, and $\beta=0.28$.
+Training is performed using the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 5e-5 for 2 epochs per dataset.
+4.2.2 Results
+As shown in Table 2, our method achieves a strong balance between accuracy and KV-Cache eviction.
+With $\alpha=1$, it maintains competitive accuracy compared to the fine-tuned Llama2-7B baseline while achieving a high mean eviction ratio of 60.00%.
+With $\alpha=0.5$, the eviction ratio decreases to 55.49%, but the average accuracy improves .
+In tasks like RTE and COPA, it even surpasses the baseline.
+This suggests that effective token eviction helps the model focus on relevant information.
+Additionally, performance varies across tasks under the same settings.
+For instance, ARC-C is more challenging to evict compared to OpenBookQA, leading to a larger accuracy drop post-eviction.
+This highlights the importance of task-specific KV-Cache eviction policies.
+Figure 3: Comparison of peak memory usage and prefilling time between the LLaMA2-7B model (without AG) and the proposed implementation (with AG and $\sim$50eviction) across varying prompt lengths.
+The results show significant improvements in memory efficiency with AG, especially as prompt length increases.
+Prefilling time is not the primary focus, and the current implementation (marked with * in the legend) relies on a suboptimal for-loop over attention heads.
+Even so, the method maintains stable prefilling time and shows a clear reduction trend with longer prompts.
+Figure 4: This visualization highlights attention patterns in Llama2-7B after fine-tuning on the BoolQ dataset, showcasing multiple heads within both MHA and AG across different layers using a selected sample.
+In part (i), we visualize attention scores from several MHA heads across layers before eviction: 1.
+MHA heads display diverse attention patterns, especially in the first two layers, where heterogeneity is prominent.
+2.
+Attention patterns become progressively sparser in deeper layers, transitioning from dense in early layers.
+3.
+Bright-yellow vertical lines, indicating critical tokens for inference, appear consistently across heads in deeper layers.
+These align with the Heavy Hitters in H2O (Zhang et al., 2024), highlighting tokens that significantly contribute to attention scores.
+Our method ensures these critical tokens are preserved in deeper layers, maintaining their importance across the network.
+In part (ii), we visualize the attention-like scores from the AG mechanism: As layers deepen, AG shifts from high-resolution to lower-resolution attention, focusing on distilling in-context information.
+Deeper AG layers no longer require high resolution for capturing global context, as earlier layers have already refined it.
+This suggests potential optimizations, such as reducing the number of heads or dimensions in deeper AG layers, to enhance efficiency.
+4.2.3 Inference Efficiency
+We evaluated the inference efficiency of the LLaMA2-7B model with and without the AG module.
+As shown in Figure 3, our method achieves notable improvements in memory efficiency, especially with longer prompts.
+Although prefilling time is influenced by the current implementation, it remains stable and demonstrates scalability as prompt lengths increase.
+We believe there is room for acceleration using kernel fusion techniques and leave it for future work.
+4.3 Visualization
+Table 3: Ablation study on various settings of AG, reporting accuracy (Acc.) and KV-Cache eviction ratio (%Evict.) under different configurations, with $\alpha=1$ in all settings.
+The results for (1) correspond to the setup described in Section 4.2, where the number of AG heads is 4, and the head dimension is 128.
+In (2-1) and (2-2), we explore the impact of the number of AG heads, with (2-1) using 2 heads and (2-2) using 1 head.
+The comparison between (1), (2-1), and (2-2) shows that reducing the number of heads leads to a drop in both accuracy and eviction ratio, indicating that the capacity of AG is closely tied to the number of heads.
+For (3-1) and (3-2), we assess the effect of reducing head dimensions for AG heads, where (3-1) has half the dimension size of (1) and (3-2) has 1/4.
+Comparing (1), (3-1), and (3-2) reveals that smaller dimensions reduce the eviction capability and accuracy, highlighting the importance of maintaining sufficient head dimensionality.
+Settings (4-1) and (4-2) explore using the previous layer’s hidden states and AG module to guide the current layer’s eviction strategy.
+In (4-1), eviction begins from the second layer onward, while in (4-2), eviction starts from the third layer, reflecting the denser attention patterns in the first two layers (Section 4.3).
+Both settings show a slight decline in accuracy and eviction ratio compared to (1) but introduce parallelism, suggesting potential for future optimization.
+The setting (5) replaces MHA in AG with a simple linear layer to determine the eviction strategy.
+The comparison with (1) shows that linear layers almost cannot evict tokens effectively, reinforcing the necessity of leveraging global in-context information for successful eviction, as discussed in Section 3.4.
+Metric PIQA ARC-C RTE COPA BoolQ OBQA Avg.
+(1) Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
+%Evict. 66.16 48.31 65.47 45.40 67.46 67.17 60.00
+(2-1) Acc. 81.88 57.63 65.70 91.00 87.52 77.40 76.86
+%Evict. 63.92 36.38 62.73 24.38 65.22 63.57 52.70
+(2-2) Acc. 82.15 53.90 62.45 89.00 87.31 77.40 75.37
+%Evict. 58.97 31.47 59.77 20.32 63.02 59.17 48.79
+(3-1) Acc. 81.45 53.36 58.84 88.00 86.73 78.40 74.46
+%Evict. 61.75 33.55 61.34 19.24 64.59 59.59 50.01
+(3-2) Acc. 83.03 53.90 59.93 89.00 87.16 76.40 74.90
+%Evict. 58.68 24.23 32.23 12.28 59.40 55.54 40.39
+(4-1) Acc. 81.66 55.25 66.06 88.00 86.85 78.00 75.97
+%Evict. 49.52 36.92 46.85 28.74 56.02 60.32 46.40
+(4-2) Acc. 82.75 55.93 79.06 82.00 86.33 78.40 77.41
+%Evict. 53.31 44.38 51.20 47.95 61.98 61.73 53.43
+(5) Acc. 82.54 54.58 57.40 81.00 87.71 74.80 73.01
+%Evict. 1.06 0.46 0.81 0.26 1.38 1.16 0.86
+In this section, we present visualizations to highlight key characteristics of our AG mechanism.
+After fine-tuning on specific tasks, we visualize the model’s MHA and AG weights for selected samples, as shown in Figure 4.
+The complete version can be found in Section A.3.
+4.4 Ablation
+In the ablation study, we investigate the effects of various configurations on the AG mechanism, including the number of AG heads, head dimensions, and eviction strategies.
+Detailed settings and performance results are provided in Table 3.
+Key findings include the impact of reducing the number of heads (as seen in (2-1) and (2-2)) and head dimensions (in (3-1) and (3-2)), both of which lead to lower accuracy and eviction ratios.
+In (4-1) and (4-2), we evaluate eviction strategies where the current layer’s eviction is based on the previous layer’s hidden states and AG module, introducing parallelism but with a slight performance trade-off.
+Replacing MHA with a linear layer in setting (5) highlights the importance of an attention-like structure for effective token eviction.
+Additional results are provided in Table 6.
+5 Conclusion
+In conclusion, the proposed Attention-Gate mechanism offers a flexible and adaptive solution to KV-Cache eviction in large language models.
+By dynamically identifying and discarding less important tokens in a data-driven manner, Attention-Gate addresses the limitations of static and attention-score-based strategies, providing efficient context-aware eviction.
+This mechanism integrates seamlessly with pre-trained models and can be easily tuned, making it a practical and effective method for enhancing both performance and memory efficiency in various tasks.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Additional Experiments
+A.1 Additional Results for Continual Pre-training
+In this section, we perform continual pre-training on Llama2-7B using the same training data and hyperparameter settings described in Section 4.1.1.
+The baselines are training-free, while in our method, only the AG module is trainable.
+The results are shown in Table 4.
 Table 4: Performance comparison of Llama2-7B and various KV-Cache eviction strategies across seven tasks.
+Our approach trains only the AG module during continual pre-training, keeping other components frozen.
 The table reports accuracy (Acc.) for Llama2-7B and all eviction methods, with Llama2-7B serving as the upper bound for accuracy.
-Metric %Eviction refers to the mean KV-Cache eviction ratio, representing the percentage of tokens evicted from the KV-Cache.
+Metric %Evict.
+refers to the mean KV-Cache eviction ratio, representing the percentage of tokens evicted from the KV-Cache.
 The eviction ratio is fixed at 50% for the baseline methods, including a local strategy (retaining only recent tokens), StreamingLLM, and H2O.
-In contrast, our method, which includes continual pre-training, achieves higher average accuracy while maintaining a higher average %Eviction.
+In contrast, our method achieves higher average accuracy while maintaining a higher average %Evict..
 Metric PIQA ARC-C ARC-E RTE COPA BoolQ HellaSwag Avg.
@@
 Ours Acc. 76.17 33.90 49.03 52.35 63.00 67.52 66.33 58.33
-%Eviction 54.29 51.03 51.05 46.70 40.02 57.75 52.16 51.87
-Table 2: Performance of Llama2-7B with LoRA fine-tuning and our method on six downstream tasks.
-In addition to the LoRA fine-tuning targets, our method makes the AG modules learnable.
-Two settings for $\alpha$ (0.5 and 1) are tested.
-Our method maintains comparable or better accuracy while achieving a higher eviction ratio, demonstrating its task-specific adaptability in managing token eviction without significant accuracy loss.
-Metric PIQA ARC-C RTE COPA BoolQ OBQA Avg.
-Fine-tuned Llama2-7B Acc. 82.92 60.34 64.98 92.00 88.10 78.80 77.86
-Ours $(\alpha=1)$ Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
-%Eviction 66.16 48.31 65.47 45.40 67.46 67.17 60.00
-Ours $(\alpha=0.5)$ Acc. 81.50 57.63 74.01 95.00 87.00 79.20 79.06
-%Eviction 64.96 36.45 62.80 34.77 67.31 66.65 55.49
-4.1.1 Setup
-Models and datasets.
-We use Llama2-7B (Touvron et al., 2023) as our primary base model due to its strong performance and popularity within the AI community.
-Additionally, we validate the feasibility of our approach on Mistral-7B (Jiang et al., 2023), with results provided in Table 4.
-For continual pre-training, we select a subset of the Redpajama dataset (Computer, 2023) containing approximately 5,000 samples  [^1] We sampled 4,997 samples proportionally from each subset of RedPajama..
-To assess the effectiveness of our method, we evaluate it on seven widely recognized benchmarks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), and HellaSwag (Zellers et al., 2019).
-All evaluations are conducted in a zero-shot setting, with accuracy assessed using OpenCompass (Contributors, 2023).
-Training details.
-During continual pre-training (CPT), only the AG modules are learnable, while the rest of the model’s parameters are frozen.
-For the Llama2-7B model, the hyperparameters from Section 3.4 are used: $\tau=0.5$, $\gamma=2$, $\alpha=5$, and $\beta=0.4$.
-The model was trained for one epoch.
-Metrics and Baselines.
-The metrics considered include accuracy and the mean eviction ratio for all KV-Cache.
-The accuracy of vanilla Llama2-7B on these datasets is used as the upper bound, with the goal of minimizing accuracy loss while maximizing the mean eviction ratio.
-For KV-Cache eviction strategies, we select a local method (retaining only recent tokens) and StreamingLLM (Xiao et al., 2024) as representatives of static strategies, while H2O (Zhang et al., 2024) serves as a representative of methods based accumulative attention scores.
-For all these methods, eviction ratio is set to 50%.
-4.1.2 Results
-The results in Table 1 demonstrate the effectiveness of our method in balancing accuracy and KV-Cache eviction.
-Our method consistently outperforms the baseline strategies in terms of accuracy across most tasks while achieving a higher mean eviction ratio.
-This indicates that our method is able to evict more tokens on average without significantly compromising accuracy.
-In terms of accuracy, our method achieves competitive performance, closely matching the performance of vanilla Llama2-7B in most cases.
-For instance, on PIQA, COPA, and BoolQ, our method performs comparably to Llama2-7B, demonstrating minimal accuracy degradation.
-Compared to baselines, our method shows superior performance, particularly in tasks like BoolQ and PIQA, where both accuracy and eviction ratio surpass the baselines.
-This confirms the advantage of learnable mechanism for more efficient token retention strategies without sacrificing model performance.
-Moreover, it is worth highlighting that our method achieves these results with minimal computational overhead, as mentioned in Section 4.1.1.
-The continual pre-training was conducted on only 5,000 samples and trained for just one epoch, demonstrating the lightweight nature of our approach.
-This efficiency can be attributed to the fact that our method does not need to learn new knowledge from scratch but rather focuses on learning effective token retention strategies, leveraging the existing capabilities of the pre-trained model.
-4.2 Supervised Fine-tuning
-Figure 3: This visualization highlights attention patterns in Llama2-7B after fine-tuning on the BoolQ dataset, using a selected sample to showcase multiple heads within both MHA and AG of different layers.
-In part (i), we visualize attention scores from several MHA heads across different layers before eviction.
-1.
-MHA heads exhibit diverse attention patterns, especially in the first two layers, where attention is highly heterogeneous across heads.
-2.
-While the first two layers show dense attention patterns, the subsequent layers become progressively sparser.
-3.
-Bright-yellow vertical lines consistently appear at the same position across nearly all heads, especially beyond the first two layers, indicating critical tokens for inference.
-This aligns with the Heavy Hitters in H2O (Zhang et al., 2024), where a small portion of tokens significantly contributes to the attention scores.
-Our method’s effectiveness is demonstrated by the persistence of these bright lines in deeper layers, showing that crucial tokens are retained rather than evicted, ensuring their importance is preserved across the network.
-In part (ii), we visualize the attention scores from the attention-like structure within the AG mechanism.
-As the model layers deepen, the attention pattern shifts from high-resolution to lower-resolution, indicating that AG increasingly focuses on distilling in-context information.
-Thus, deeper AG layers no longer need such high resolution to capture global information, as the model has already refined and summarized it.
-This suggests potential optimizations, such as reducing the number of heads or dimensionality in deeper AG layers, to further improve efficiency.
-4.2.1 Setup
-Model and tasks.
-We utilize Llama2-7B (Touvron et al., 2023) as base model.
-To evaluate our approach, we select six widely recognized downstream tasks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), and OpenBookQA (Mihaylov et al., 2018).
-For each task, we fine-tune the model using the respective training set and evaluate its performance on the corresponding test set.
-Baselines and Implementation Details.
-For the baseline, we apply LoRA (Hu et al., 2021) to fine-tune the models on each task, targeting the $(W_{q},W_{k},W_{v},W_{o})$ weights in the self-attention modules.
-In our method, in addition to applying LoRA to these parameters, the AG modules are also set to be learnable.
-As mentioned in Section 3.4, the following hyperparameters are used: $\tau=0.3$, $\gamma=0$, $\alpha=1$ or $0.5$, and $\beta=0.28$.
-We employ the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 5e-5, and train for two epochs on each dataset.
-4.2.2 Results
-As shown in Table 2, our method demonstrates a strong balance between accuracy and KV-Cache eviction across the six downstream tasks.
-With $\alpha=1$, our method maintains competitive accuracy compared to the fine-tuned Llama2-7B baseline, with minimal drops (e.g., 82.15% vs.
-82.92% on PIQA), while achieving a high mean eviction ratio of 60.00%.
-With $\alpha=0.5$, the eviction rate is reduced to 55.49%.
-But in some tasks like RTE and COPA, our method even surpasses the baseline accuracy.
-The average accuracy also exceeds the baseline (79.06% vs.
-77.86%), suggesting that effectively evicting redundant tokens allows the model to focus on relevant information and improves performance.
-Additionally, under the same hyperparameter settings, the performance varies across tasks.
-For instance, ARC-C is more challenging to evict compared to OpenBookQA, leading to a larger accuracy drop post-eviction.
-This highlights the importance of task-specific KV-Cache eviction policies.
-4.3 Visualization
-Table 3: Ablation study on various settings of AG, reporting accuracy (Acc.) and KV-Cache eviction ratio (%Eviction) under different configurations, with $\alpha=1$ in all settings.
-The results for (1) correspond to the setup described in Section 4.2, where the number of AG heads is 4, and the head dimension is 128.
-In (2-1) and (2-2), we explore the impact of the number of AG heads, with (2-1) using 2 heads and (2-2) using 1 head.
-The comparison between (1), (2-1), and (2-2) shows that reducing the number of heads leads to a drop in both accuracy and eviction ratio, indicating that the capacity of AG is closely tied to the number of heads.
-For (3-1) and (3-2), we assess the effect of reducing head dimensions for AG heads, where (3-1) has half the dimension size of (1) and (3-2) has 1/4.
-Comparing (1), (3-1), and (3-2) reveals that smaller dimensions reduce the eviction capability and accuracy, highlighting the importance of maintaining sufficient head dimensionality.
-Settings (4-1) and (4-2) examine using the hidden states and the AG module from the previous layer to inform the current layer’s eviction strategy.
-In (4-1), the first layer does not evict, and from the second layer onward, the eviction is determined by the previous layer.
-In (4-2), the first and second layers do not evict, following the observation from Section 4.3 that the first two layers are denser than subsequent layers.
-Both (4-1) and (4-2) show a slight decline in accuracy and eviction ratio compared to (1), but this approach introduces parallelism, offering a potential avenue for future optimizations.
-The setting (5) replaces the attention-like structure in AG with a simple linear layer to determine the eviction strategy.
-The comparison with (1) shows that linear layers almost cannot evict tokens effectively, reinforcing the necessity of leveraging global in-context information for successful eviction, as discussed in Section 3.3.
-Metric PIQA ARC-C RTE COPA BoolQ OpenBookQA Avg.
-(1) Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
-%Eviction 66.16 48.31 65.47 45.40 67.46 67.17 60.00
-(2-1) Acc. 81.88 57.63 65.70 91.00 87.52 77.40 76.86
-%Eviction 63.92 36.38 62.73 24.38 65.22 63.57 52.70
-(2-2) Acc. 82.15 53.90 62.45 89.00 87.31 77.40 75.37
-%Eviction 58.97 31.47 59.77 20.32 63.02 59.17 48.79
-(3-1) Acc. 81.45 53.36 58.84 88.00 86.73 78.40 74.46
-%Eviction 61.75 33.55 61.34 19.24 64.59 59.59 50.01
-(3-2) Acc. 83.03 53.90 59.93 89.00 87.16 76.40 74.90
-%Eviction 58.68 24.23 32.23 12.28 59.40 55.54 40.39
-(4-1) Acc. 81.66 55.25 66.06 88.00 86.85 78.00 75.97
-%Eviction 49.52 36.92 46.85 28.74 56.02 60.32 46.40
-(4-2) Acc. 82.75 55.93 79.06 82.00 86.33 78.40 77.41
-%Eviction 53.31 44.38 51.20 47.95 61.98 61.73 53.43
-(5) Acc. 82.54 54.58 57.40 81.00 87.71 74.80 73.01
-%Eviction 1.06 0.46 0.81 0.26 1.38 1.16 0.86
-In this section, we present visualizations to highlight key characteristics of our AG mechanism.
-After fine-tuning on specific tasks, we visualize the model’s MHA and AG weights for selected samples, as shown in Figure 3.
-Additional visualizations can be found in Section A.2.
-4.4 Ablation
-In this ablation study, we explore the effects of various configurations on the AG mechanism, focusing on the number of AG heads, AG head dimensions, and eviction strategies.
-For details of each setting and the corresponding performance, refer to Table 3.
-Key findings include the impact of reducing the number of heads, as shown in (2-1) and (2-2), and the reduction of head dimensions in (3-1) and (3-2), both of which result in lower accuracy and eviction ratios.
-In (4-1) and (4-2), we examine eviction strategies where the current layer’s eviction is based on the previous layer’s hidden states and AG module, which introduces parallelism but slightly reduces performance.
-Finally, replacing the AG with a linear layer in (5) demonstrates the necessity of using an attention-like structure for effective token eviction.
-For more ablation, please refer to Table 5.
-5 Conclusion
-In conclusion, the proposed Attention-Gate mechanism offers a flexible and adaptive solution to KV-Cache eviction in large language models.
-By dynamically identifying and discarding less important tokens, Attention-Gate addresses the limitations of static and attention-score-based strategies, providing efficient context-aware eviction.
-This mechanism integrates seamlessly with pre-trained models and can be easily tuned, making it a practical and effective method for enhancing both performance and memory efficiency in various tasks.
-Appendix A Appendix
-A.1 Results of Continual Pre-training on Mistral
-We conducted continual pre-training on Mistral using 5,000 samples from RedPajama (Computer, 2023), and the results are shown in Table 4.
+%Evict. 54.29 51.03 51.05 46.70 40.02 57.75 52.16 51.87
+A.2 Results of Continual Pre-training on Mistral
+We conducted continual pre-training on Mistral-7B (Jiang et al., 2023) using 5,000 samples from RedPajama (Computer, 2023), and the results are shown in Table 5.
 Compared to the performance of Llama2-7B presented in Table 4, Mistral’s performance slightly declined.
@@
 This could suggest that Mistral is utilizing more tokens or scoring them with finer granularity, which results in fewer redundant tokens and thus makes eviction less effective.
+Furthermore, it is possible that Mistral’s use of grouped-query attention (GQA), which inherently involves compression, may make it more challenging to increase the eviction ratio effectively in this context.
 Table 5: Performance comparison between Mistral-7B and Ours across various tasks.
@@
 Mistral-7B Acc. 80.09 42.37 63.14 48.01 76 64.22 73.02 63.84
-Ours Acc. 75.9 34.24 55.2 48.01 65 62.2 67.91 58.35
+Ours Acc. 75.90 34.24 55.2 48.01 65 62.2 67.91 58.35
 Eviction 37.14 39.48 37.80 40.93 45.27 44.68 50.92 42.32
-A.2 More Visualization
+A.3 More Visualization
+Figure 5 provides a comprehensive view of the layers and attention heads from Figure 4.
 Figure 5: The complete version of Figure 4.
-Figure 5: More samples for the same scenario of Figure 3.
-Figure 4 provides a comprehensive view of the layers and attention heads from Figure 3. Additionally, four sample scenarios for the same setup are presented in Figure 5.
-A.3 More Ablation
+Two key observations emerge in (i): 1.
+the first two layers are denser compared to the subsequent layers, and 2.
+bright-yellow vertical lines, representing critical tokens for inference, consistently appear across heads in deeper layers.
+A.4 More Ablation
+In this section, we present additional ablation results in Table 6.
 Table 6: Exploring the impact of the number of recent tokens, viewed from the perspective of the attention matrix and considering slanted retention patterns.
@@
 (1) Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
-%Eviction 66.16 48.31 65.47 45.40 67.46 67.17 60.00
+%Evict. 66.16 48.31 65.47 45.40 67.46 67.17 60.00
 (6-1) Acc. 83.08 50.85 65.34 82.00 87.31 73.20 73.63
-%Eviction 65.15 40.96 64.29 21.37 67.49 63.69 53.83
+%Evict. 65.15 40.96 64.29 21.37 67.49 63.69 53.83
 (6-2) Acc. 81.61 53.56 60.29 82.00 87.37 74.20 73.17
-%Eviction 65.66 44.48 65.14 24.28 68.18 63.44 55.20
-In this section, we present additional ablation results in Table 5.
+%Evict. 65.66 44.48 65.14 24.28 68.18 63.44 55.20